If bock 1 weighs 50 kg, while block 2 weighs 90 kg , determine the external force p2 so that block 2 will maintain its position. Let p1 be equal to 5kN.Assume 0.3 coefficient of statics friction




1
Expert's answer
2021-11-07T19:26:42-0500

Block 1 is on top of block 2. For the block to stay at rest, the total force "p_1+p_2" must not exceed the force of static friction:


"f_s=\mu_sN=\mu_sg(m+M),\\\space\\
-f_s+p_1+p_2=0,\\
p_2=f_s-p_1=\mu_sg(m+M)-p_1=-4588.4\text{ N}"

opposite to p1.